OpenStudy (anonymous):

f(x)=e^x+x and its inverse is g(x). Calculate g'(1)

OpenStudy (anonymous):

have you found the inverse?

OpenStudy (anonymous):

I know d/dx e^x is e^x.

OpenStudy (anonymous):

the inverse is just 1/(e^x+x) i beleive

OpenStudy (anonymous):

\[f(x)=e^{x}+x\] is your equation right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

the inverse of differentiation is integration, or finding the antiderivative anti derivative of x is x^2/2, so that when you differentiate it you get x back

OpenStudy (anonymous):

problem is because we have not covered that part in class yet we cant use in in our solution.

OpenStudy (anonymous):

that and I have no idea what you mean by the antiderivative

OpenStudy (anonymous):

so in fact you dont need to do any integration here, pay close attetion to the question g'(x) = f(x) so all you need to do is eval f(1) to get g'(1)

OpenStudy (anonymous):

hmm then just throw it over 1, but in this case it will still be 1

OpenStudy (anonymous):

g'(1)=f(1)=e+1

OpenStudy (anonymous):

well here let me scan the question as thats what I did but apparently its wrong

OpenStudy (anonymous):

@k1dj03y it states that g(x) is the inverse not the derivative

OpenStudy (anonymous):

f(x)=e^x+x and its inverse g(x). Calculate g'(1)

OpenStudy (anonymous):

oops i added the word is to that question.

OpenStudy (anonymous):

it's still the same thing, the question tells you that the inverse of f(x) is g(x)

OpenStudy (anonymous):

so if you differentiate the inverse function you will get the original function back, which is f(x)

OpenStudy (anonymous):

hmm thats just it then as when it was graded i only received 3/10 possible for the solution.

OpenStudy (anonymous):

@arjont you are thinking of intergration that is not what the relationship between f(x) and g(x) is

OpenStudy (anonymous):

Im attaching the question and what I received back as feedback to see if it helps things

OpenStudy (anonymous):

The graph may make things look better as he gives us g(x) in the form of a graph. Which also happens to be the inverse of (f(x)) if I am understanding things right

OpenStudy (anonymous):

I could just look at the graph and be like there it is, but, it asks for it to be calculated. . .

OpenStudy (anonymous):

I think this is the formula your teacher wanted you to use \[[f ^{-1}]\prime(a)=\frac{ 1 }{ f \prime(f ^{-1}(x)) }\] look familiar?

OpenStudy (anonymous):

somewhat yeah

OpenStudy (anonymous):

well you can solve for g(1) without finding the inverse function. You can also calculate the derivative of f(x) plug them into that equation and you should get you answer

OpenStudy (anonymous):

@k1dj03y I made a mistake typing my question that x variable should be a, sorry about that

OpenStudy (anonymous):

yeah we have been using u but its the same idea so i understood where it was going at least

OpenStudy (anonymous):

do you need more help are do you think you are good with this question now?

OpenStudy (anonymous):

naw I think I can take it from there thx bunches

OpenStudy (anonymous):

Okay don't forget to close the question.

OpenStudy (anonymous):

haha yeah I was going to. Just hate cutting people off mid typing

OpenStudy (anonymous):

@arjont do you want/need further explanation?

OpenStudy (anonymous):

no, i didnt know what he meant by g' i thought it was the derivative

OpenStudy (anonymous):

they do want you to find the value of the derivative of the inverse at one. Its just an error in reading the question then.